### v1

# Accelerating Speculative Decoding using Dynamic Speculation Length

Abstract
Speculative decoding is a promising method for reducing the inference latency of large language models. The effectiveness of the method depends on the speculation length (SL)—the number of tokens generated by the draft model at each iteration. The vast majority of speculative decoding approaches use the same SL for all iterations. In this work, we show that this practice is suboptimal. We introduce DISCO, a DynamIc SpeCulation length Optimization method that uses a classifier to dynamically adjust the SL at each iteration, while provably preserving the decoding quality. Experiments with four benchmarks demonstrate average speedup gains of 10.3% relative to our best baselines. [^1] Code will be available upon publication.
Accelerating Speculative Decoding using Dynamic Speculation Length
Jonathan Mamou  Oren Pereg  Daniel Korat Moshe BerchanskyNadav TimorMoshe WasserblatRoy Schwartz Intel Labs, Israel Weizmann Institute of Science, Israel School of Computer Science & Engineering, Hebrew University of Jerusalem Correspondence: jonathan.mamou@intel.com
1 Introduction
Large language models (LLMs) generate tokens in an autoregressive manner, incurring slow and costly generation. Speculative decoding methods Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023) mitigate this issue by splitting the inference of a given target model into two steps. In the first step, a faster draft model generates tokens autoregressively. In the second step, the target model validates all the draft tokens simultaneously. The validation procedure accepts or rejects the draft tokens (Fig. 1). This approach accelerates the generation process by producing multiple tokens in a single forward pass of the target model, while maintaining the decoding quality of the target model.
The effectiveness of speculative decoding in expediting the token generation process depends on the speculation length (SL) parameter, which pertains to the number of tokens generated by the draft model before each validation step (Chen et al., 2023). The choice of SL has a large effect on latency: lower static (constant) SL values may incur a lower number of accepted draft tokens per speculative iteration, which leads to more target model validations than necessary. In contrast, higher static SL may result in a large number of ignored draft generations in case of an early draft-target mismatch (Fig. 1).
Figure 1: An illustration of a single speculative decoding iteration with static speculation length ($SL$) $=5$. A draft model generates $SL$ tokens autoregressively. The target model validates them all in parallel, and accepts the first three tokens. In this case, $SL$ is too high (as the last two draft tokens are never used). $SL$ can also be too low (e.g., $SL=2$ here), which would lead to too many calls to the target model.
The vast majority of speculative decoding approaches Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023); Cai et al. (2024); Li et al. (2024) employ a static SL, i.e., a constant number of draft tokens are generated in each speculative iteration. In this work, we show that this approach is suboptimal. We start by introducing a dynamic oracle SL, which sets the SL to be the exact number of draft generations until the first mismatch in each speculative iteration. We observe a high variance in the optimal SL across the speculative iterations. This finding demonstrates the potential of dynamically adapting the speculation length.
We then propose DISCO, a dynamic speculation length optimization method. DISCO uses draft model features to train a lightweight classifier that predicts the compatibility between a generated draft token and the target token. At inference time, the classifier determines whether the draft model should continue generating the next token, or halt and switch to the target model for validation.
We evaluate DISCO’s effectiveness in code generation, text summarization and instruction tasks. Our results demonstrate average speedup gains of 10.3% and 31.4% relatively to the best static and dynamic SL baselines, respectively, while maintaining the same generated text. Our classifier also transfers well across tasks from the same category: training it on one task and using it on another leads to similar gains over our baselines.
Figure 2: The improvement factor (IF) as a function of the speculation length ($\gamma$) for three different acceptance rate ($\alpha$) values assuming fixed cost coefficient ($c=0.05$). The IF value largely depends on both $\alpha$ and $\gamma$.
2 Background: Speculative Decoding
Speculative decoding Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023) expedites LLM generation while ensuring no accuracy loss by dividing it into two stages. In the first stage, a fast but less accurate draft model $M_{D}$ autoregressively generates a sequence of tokens. In the second stage, a large but more accurate target model $M_{T}$ conducts parallelized verification over the generated draft tokens. This process allows the model to potentially produce multiple tokens per target forward pass.
Speculation length (SL)
The effectiveness of speculative decoding in accelerating the token generation process relies heavily on the SL parameter, which determines how many tokens are generated by the draft model before each validation step. The effect of the SL on the overall speedup is subject to a tradeoff; higher SLs potentially reduce the number of target model validations, but also increase the number of redundant draft generations (Fig. 1).
The majority of speculative decoding approaches Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023); Cai et al. (2024); Li et al. (2024) use a static SL—the same number of draft tokens are generated per speculative iteration.222 Some practical work apply heuristics for dynamic SL adaptation, increasing/decreasing SL when all/some draft tokens are accepted/rejected in the previous iteration (https://huggingface.co/blog/assisted-generation). Chen et al. (2023) explored various static SLs, across different target-draft model pairs and tasks, and empirically showed that as the SL rises, the overall speedup increases until reaching a certain threshold, beyond which it either levels off or even regresses.
To study the effect of different static SL values, Leviathan et al. (2023) defined the improvement factor (IF) as the expected token generation speedup. Formally:
$IF={\frac{1-\alpha^{\gamma+1}}{(1-\alpha)({\gamma}c+1)}}$ (1)
where $\alpha$ denotes the acceptance rate, indicating the expected probability of a draft token to be accepted by the target model; $c$ represents the cost coefficient, indicating the ratio between the walltime of a forward pass run of the draft model $M_{D}$ and the wall time of a forward pass run of the target model $M_{T}$; and $\gamma$ represents the static SL value. [^3] IF computation assumes enough compute resources for increased concurrency as $\gamma$ rises.
Figure 2 shows the IF as a function of $\gamma$ for three different values of $\alpha$ and a fixed $c$ value. While both $\alpha$ and $c$ are important to the selection of the target-draft model pair, finding the optimal $\gamma$ is fundamental to the effectiveness of the system.
3 Dynamic Speculation Length
The IF function is based on the simplifying assumption that the probability of accepting draft tokens by the target model is i.i.d. Nevertheless, in practical scenarios, different tokens may have varying levels of predictability, rendering the simplifying i.i.d. assumption unsound. Therefore, relying on a static SL might not be optimal, as certain speculative iterations may benefit from generating more tokens than others. Below we consider an oracle experiment, which applies the optimal dynamic $\gamma$ value at each iteration. We then propose a method for dynamically setting $\gamma$, showing that it strongly outperforms a static selection method for any choice of a static SL.
Assessing the optimal dynamic $\gamma$ value
We start by employing an oracle for detecting the ‘ground truth’ value of $\gamma$ for each speculative iteration. The oracle uses the draft model to autoregressively generate tokens until a mismatch occurs between the predicted tokens of the draft and target models. This process is repeated for each speculative iteration, ultimately returning the optimal (maximum) number of accepted draft tokens per iteration. The mismatch between the tokens is determined by using the rejection sampling algorithm introduced by Leviathan et al. (2023) with zero temperature. This oracle fulfills the speculative decoding potential: generating the maximal number of valid draft tokens at each iteration, and making a minimal number of calls to both draft and target model.
Figure 3 shows the oracle SL values across the speculative iterations for one MBPP example. Comparing with static SL, we observe a lower number of oracle draft and target forward passes. Figure 4 shows the average oracle SL over the speculative iterations for Alpaca dataset. We observe on both figures a high variance of oracle SL values, implying that setting a predefined static SL is likely to be suboptimal. See App. A for further analysis.
Figure 3: Oracle and static SL values for different speculative iterations on one MBPP example. For static SL, we run 38 target forward passes and 192 draft forward passes, while for oracle SL, we only run 27 target forward passes and 129 draft forward passes. We observe a high variance of oracle SL values.
Figure 4: Bar chart of the average oracle SL over the normalized index of the speculative iterations for the Alpaca dataset. We observe a high variance of oracle SL values.
DynamIc SpeCulation length Optimization
We introduce DISCO, a simple method for dynamically setting the SL value at each iteration. To estimate the correct SL value at each step, we employ a simple SL classifier. Our classifier decides, at each draft generation step, whether the draft model $M_{D}$ should proceed to generate the next token or switch to the target model $M_{T}$ for verification. The classifier takes as inputs the probability vector of the draft model ($y^{D}_{i}$) and the token position ($i$), generating a confidence score ($C_{i}$) used for the decision-making as follows:
$C_{i}=FFN(Concat(Top_{k}(y^{D}_{i}),Ent(y^{D}_{i}),i))$
where $Top_{k}()$ selects the top $k$ values [^4] We use $k=10$ in all experiments. and $Ent()$ is the entropy function.
[…]
4 Experiments
Datasets
We evaluate our method on four datasets spanning three tasks: code generation, using MBPP Austin et al. (2021) and HumanEval Chen et al. (2021); text summarization using CNN-DailyMail Nallapati et al. (2016); and instruction-following using Alpaca Taori et al. (2023). We use the training set for training the SL classifier and the validation set for setting threshold $\tau$ and $SL_{max}$ hyperparameters. For HumanEval, which has no training and validation sets, we evaluate transfer learning from MBPP. See App. B for more details on the datasets and prompts.
Models
To evaluate speculative decoding on code generation tasks, we use the Starcoder model family Li et al. (2023)—15B for target and 168M for draft. For our text generation tasks, we use Vicuna models—13B as target Chiang et al. (2023) and 68M as draft Yang et al. (2024). Throughout the experiments we use the rejection sampling scheme Leviathan et al. (2023) with the temperature set to zero. Exact model names and versions are provided in App. C.
SL classifier training
For minimal overhead, we use a shallow 2-layer FFN classifier. We train the SL classifier to predict the conformity between $M_{D}$ and $M_{T}$ by employing cross-entropy loss with the Total Variance distance: $TV(y^{D}_{i},y^{T}_{i})$, where $y^{D}_{i}$ and $y^{T}_{i}$ represent the vocabulary distribution of $M_{D}$ and $M_{T}$ respectively at the $i^{th}$ token position. We evaluate the quality of the classifier by measuring its F1 score on the validation set. The F1 results obtained on the datasets are relatively high; for instance, the F1 score on the MBPP dataset is 95% while the F1 score of the optimal static SL is 85%. For further details, see App. E and App. D.
Baselines
We compare the LLM inference latency of DISCO to both static SL (Static SL-5) and dynamic heuristic SL setups (dynHeur SL; see Fn. 2). We also consider the optimal static SL baseline tuned on our validation sets (Static SL-opt). Finally, we compare DISCO to our oracle (Sec. 2) that represents the lower bound on latency.
Results
Dataset Method Latency(ms) Speedup
MBPP Target 23.21 1.00x
dynHeur SL 20.07 1.16x
Static SL-5 15.88 1.46x
Static SL-opt 14.16 1.64x
DISCO (ours) 12.58 1.84x
Oracle 10.18 2.28x
HumanEval* Target 23.46 1.00x
dynHeur SL 22.57 1.04x
Static SL-5 14.43 1.63x
Static SL-opt 14.42 1.63x
DISCO (ours) 12.78 1.84x
Oracle 10.59 2.22x
CNN-DM Target 38.29 1.00x
dynHeur SL 21.18 1.81x
Static SL-5 19.74 1.94x
Static SL-opt 20.66 1.85x
DISCO (ours) 17.85 2.15x
Oracle 15.41 2.48x
Alpaca Target 47.67 1.00x
dynHeur SL 31.83 1.50x
Static SL-5 23.79 2.00x
Static SL-opt 23.65 2.02x
DISCO (ours) 22.49 2.12x
Oracle 20.04 2.38x
Table 1: Average latency results on different benchmarks. HumanEval results (marked with *) use a classifier trained on MBPP (transfer learning).
Table 1 presents our results. Employing an SL classifier consistently outperforms all other baselines across all benchmarks. Average latency improvements of DISCO over the optimal static SL and the dynamic heuristic baselines are 10.3% and 31.4% respectively (see App. F), while preserving the same output as the target model. Importantly, our improvement does not come only from our training data: the optimal static SL (as fit by that data) is still underperformed by DISCO. Finally, we note that classifiers transfer well across tasks: when trained on MBPP, DISCO is still able to outperform all baselines on HumanEval.
5 Related Work
There has been a growing body of work in the emerging field of speculative decoding for LLM acceleration. Pioneering studies by Leviathan et al. (2023) and Chen et al. (2023) introduced a novel rejection sampling scheme which preserves the distribution of the target model, effectively guaranteeing that speculative decoding setups will maintain the quality of the target model. Subsequent work Miao et al. (2024) elevated the average number of accepted tokens by using several draft models. More recent works Cai et al. (2024); Li et al. (2024); Bhendawade et al. (2024) eliminated the need for a separate draft model by incorporating specialized draft layers on top of the transformer decoder to forecast a sequence of draft tokens. All of the above works employed a static SL.
A very recent concurrent work by S et al. (2024) enhanced the draft model’s accuracy by granting it access to the target model’s representations. In addition, it employed a classifier to determine whether to halt or continue the speculation process. In this sense, this work is the closest to ours; however, our work delves into the impact of the SL on the efficiency of speculative decoding, encompassing comparisons between static and dynamic SL approaches, as well as the upper bound of improvement represented by the oracle SL.
6 Conclusion
We introduced DISCO, a dynamic speculation length optimization method. The method uses a classifier that determines whether the draft model should continue to generate the next token or pause and transition to the target model for validation. We evaluated DISCO’s effectiveness using four benchmarks and demonstrated average speedup gains of 10.3% and 31.4% relatively to the optimal static SL and dynamic heuristic baselines. Our results highlight the potential of further reducing inference cost by using simple, efficient techniques.
7 Limitations
While our study shows improved effectiveness of speculative decoding dynamic SL adaptation, several limitations should be acknowledged.
First, although we conducted experiments on four different datasets, our evaluation was limited to only two pairs of models, namely Vicuna and Starcoder model families. Extending the analysis to a broader range of model configurations could provide a more comprehensive understanding of the generalizability of our findings.
Second, we note that the effectiveness of dynamic speculation length (SL) approaches is contingent upon the standard deviation of the SL as measured by the oracle. When the standard deviation is low, there may be less room for improvement using dynamic SL strategies.
8 Potential Risks
As this work is dedicated to evaluating existing models on publicly available datasets, we are not aware of any potential negative impact.
Appendix A Oracle SL Analysis
Table 2 shows the oracle SL expectancy and standard deviation of the oracle measured on different datasets and models and Figure 5 shows the probability distribution of the oracle SL for the different datasets. We observe a high variance of SL values.
Datasets Target Model Draft Model Oracle SL MBPP Starcoder-15B Starcoder-168M 18.0 ($\pm$ 46.2) HumanEval Starcoder-15B Starcoder-168M 14.7 ($\pm$ 42.6) CNN-DM Vicuna-13B Vicuna-68M 03.2 ($\pm$04.0) Alpaca Vicuna-13B Vicuna-68M 02.4 ($\pm$02.1)
Table 2: Average (and STD) of the oracle SL as determined by the oracle per dataset and target/draft models.
Figure 5: Oracle SL probability histogram on the different datasets. We observe a high variance of SL values.
Fig. 6 shows bar charts of the average oracle SL over the normalized index of the Speculation Iteration. We calculate the bars as follows. For each prompt of a dataset, we have its corresponding sequence of oracle SLs. The length of the sequence is equal to the number of Speculation Iterations. For example, consider a prompt with oracle SLs $\langle 7,3,13,21,8\rangle$. Its normalized index is $\langle 0,0.25,0.5,0.75,1\rangle$. The bars represent the average oracle SL of buckets of size $0.0001$.
(a) MBPP
(b) HumanEval
(c) CNN-DM
(d) Alpaca
Figure 6: Bar chart of the average oracle SL. Note that the iteration index seems to have low predictive power for the oracle SL.
Appendix B Datasets and Prompts Details
We use standard datasets from Hugging Face and standard prompts from the state-of-of-the-art. Tab. 3 summarizes the composition of the datasets. We provide more details per dataset in the next sections.
Datasets Train Validation Test
MBPP 374 80 80
HumanEval - - 80
CNN-DM 500 80 80
Alpaca 500 80 80
Table 3: Number of samples per dataset and split.
B.1 MBPP
For MBPP, we use the ‘train’, ‘validation’ and ‘test’ splits of the ‘full’ subset. The whole ‘train’ split is used for training, while 80 randomly selected samples of the ‘validation’ and ‘test’ splits are respectively used for validation and test. MBPP is distributed under the cc-by-4.0 License.
Concerning the prompt, we followed Ben Allal et al. (2022); Fried et al. (2023) and included the description of the programming task and a single test to verify solution, in order to help the model catch the signature of the function (see Fig. 7).
"""{text}
{test_list[0]}
Figure 7: MBPP Prompt
B.2 HumanEval
HumanEval dataset contains a single subset with a single split (‘test’ split). We use 80 randomly selected samples of that split for test. Note that since we evaluate transfer learning from MBPP, we don’t need HumanEval training and validation sets. HumanEval is distributed under the MIT License.
Prompt contains only prompt field from the dataset.
B.3 CNN-DM
For CNN-DM, we use the ‘train’, ‘validation’ and ‘test’ splits of the ‘2.0.0’ subset. 500 randomly selected samples of the ‘train’ split is used for training, while 80 randomly selected samples of the ‘validation’ and ‘test’ splits are respectively used for validation and test. CNN-DM is distributed under the Apache License 2.0.
We included the article field in the prompt as in Fig. 8.
"""Summarize:
{article}
Summary:
Figure 8: CNN-DM Prompt
B.4 Alpaca
Alpaca dataset contains a single split (‘train’ split). As for CNN-DM, 500 randomly selected samples of the ‘train’ split is used for training, while 80 randomly selected samples of the ‘validation’ and ‘test’ splits are respectively used for validation and test. Alpaca is distributed under the cc-by-nc-4.0 License.
We follow Taori et al. (2023) to define the prompts. For samples with a non-empty input field, we use the prompt as in Fig. 9 while for samples with empty input field, we use the prompt as in Fig. 10.
 """Below is an instruction that describes a
 task, paired with an input that provides
 further context. Write a response that
 appropriately completes the request.
### Instruction:
{instruction}
### Input:
{input}
### Response:
Figure 9: Alpaca prompt for samples with a non-empty input field.
"""Below is an instruction that describes a
task. Write a response that appropriately
completes the request.
### Instruction:
{instruction}
### Response:
Figure 10: Alpaca prompt for samples with empty input field.
Appendix C Models
For all models, we retrieve model weights from Hugging Face. For clarity and reproducibility, we provide the URLs for each model used:
Appendix D Classifier
D.1 Classifier Training Data
In order to train the model we created training data using the following method: the target model followed the ‘standard’ autoregressive approach to generate tokens based on the input prompt, whereas the draft model iteratively generated a single token per iteration. In each iteration, the draft model generates the next token based on the concatenation of the input prompt and the tokens subsequently generated by the target model. Draft tokens resembling the target token at the same position were labeled ‘accepted’, while others were labeled ‘rejected’. Corresponding features were extracted for both draft and target tokens, encompassing the top-k probabilities of the vocabulary distribution, the entropy associated with these probabilities, and the token position value counted from the beginning of the generation process.
D.2 Classifier F1 Results
We evaluate the quality of the classifier by measuring its F1 score on the validation set and report in Table 4 F1 scores for both static SL and DISCO. F1 scores of DISCO always outperforms F1 scores of static SL.
Datasets F1 static-SL F1 DISCO
MBPP 85 95
CNN-DM 76 88
Alpaca 68 77
Table 4: Classifier F1 scores for both static SL and DISCO.
Appendix E Additional Implementation Details
Our implementation is based on the Transformers library of HuggingFace, distributed under the Apache License 2.0, and PyTorch Deep Learning library, distributed under the BSD License (BSD-3). Our code will be available upon publication under the Apache License 2.0.
For every dataset, DISCO classifier is trained on the train set; threshold $\tau$ and $SL_{max}$ hyper-parameters are fine-tuned on the validation set optimizing the latency. Optimal static SL is estimated on the validation set. Latency results are reported on the test set.
All our experiments are run on a single A100 80GB GPU.
Appendix F Additional Results
Table 5 shows the percentage of improvement in latency of DISCO over Static SL-opt and dynHeur SL baselines. The numbers are calculated based on the latency results shown in Table 1
Dataset Static SL-opt dynHeur SL
MBPP 11.2 37.3
HumanEval 11.4 43.4
CNN-DM 13.6 15.7
Alpaca 4.9 29.3
Average 10.3 31.4
Table 5: The latency improvement (percentage) of DISCO over Static SL-opt and dynHeur SL across four datasets.